# Evaluating the Evaluators: Metrics for Compositional Text-to-Image Generation

### v1

@@ 1 Introduction
Recent advances in multi-modal generative models, such as Stable Diffusion [26, 24, 6] and DALL-E [25], have made it possible to generate high-quality, natural, and diverse images from textual descriptions at low cost and on a large scale. Alongside this progress, a parallel line of research has emerged around a central question: how can we faithfully assess the alignment between text and image, particularly in terms of compositional alignment?
Text–image compositional alignment refers to how well the objects, attributes, and relations described in a prompt are reflected in the generated image. A variety of metrics have been introduced to evaluate this alignment. Content-based metrics, such as VQAScore [20], TIFA [11], DA Score [29], DSG [5], and B-VQA [12], assess compositional properties through structured queries to the image. Embedding-based metrics, including CLIPScore [10], PickScore [16], HPS [31], and ImageReward [32], measure alignment via text–image similarity in a shared representation space or by using models trained on human preference data. Image-only metrics, such as CLIP-IQA [30] and Aesthetic Score [28], instead focus on perceptual quality and visual appeal, complementing alignment-based evaluation. Together, these metrics determine how models are evaluated and compared, and much of the field’s reported progress is defined through their outcomes. Selecting appropriate metrics is therefore essential.
In addition to their role as evaluation tools, text–image alignment metrics have increasingly served as reward signals to enhance compositional generation in diffusion models through both inference-time and reinforcement learning strategies. For instance, ReNO [7] leverages multiple reward models, including CLIPScore [10], ImageReward [32], PickScore [16], and HPSv2 [31], to guide gradient-based optimization of initial noise during inference, improving both faithfulness and aesthetics. In contrast, ImageSelect [15] adopts a best-of-$N$ sampling strategy, generating multiple candidate images from different initial noises and selecting the one with the highest ImageReward [32] score. On the training side, reinforcement learning methods like DPOK [8] fine-tune the diffusion model by maximizing ImageReward [32], using online policy gradient to strengthen compositional alignment.
Despite their widespread use in both evaluation and generation, there has been no thorough and comparative analysis of how well these metrics reflect human judgment. To address this gap, this work presents a comprehensive evaluation of $12$ text–image compositional alignment metrics on $2400$ generated text–image samples spanning $8$ compositional categories [13], examining how closely each metric corresponds to human assessments.
@@ 2.1 Evaluation Metrics
Embedding-based metrics evaluate alignment by comparing text–image representations in a shared multimodal space or by leveraging models trained on human preferences. A common baseline is CLIPScore [10], which measures cosine similarity between CLIP embeddings. Preference-supervised variants include HPS [31], which fine-tunes CLIP on human comparisons, and PickScore [16], which learns from pairwise preference judgments. BLIP [18] follows the embedding-similarity approach, comparing captions generated from images with the input text. Extending this idea, ImageReward [32] adds a reward head trained on ranked human preference data, capturing both textual relevance and perceptual quality.
@@
VQA-based metrics assess compositional alignment by casting text–image consistency as a question answering task. Questions derived from the prompt are posed to a pretrained VQA model, with scores based on the correctness of its responses. VQAScore [20] generates yes/no questions from the text, while TIFA [11] uses structured templates to cover objects, attributes, and relations. Variants target specific aspects: DA Score [29] asks entity–attribute questions to test binding, DSG [5] converts the text into a scene graph to verify entities and relations, and B-VQA [12] decomposes the text into object–attribute pairs, querying each with BLIP-VQA and combining the probabilities.
@@
Image-only metrics assess perceptual quality independently of the prompt, providing complementary signals of realism and aesthetics. CLIP-IQA [30] predicts image quality by regressing CLIP embeddings against human quality annotations, while the Aesthetic Score [28] estimates aesthetic value from large-scale human ratings.
@@ 2.2 Benchmarks
Text–image compositional alignment benchmarks, such as T2I-CompBench++ [13], group prompts into categories that reflect major alignment challenges. These include: entity existence (all entities appear without omission or hallucination), attribute binding (objects match specified properties like color or shape), spatial relations (2D and 3D object arrangements, e.g., “a cube on a sphere” or “a ball inside a box”), non-spatial relations (functional interactions, e.g., “a man holding a guitar”), and numeracy (accurate object counts, e.g., “five birds on a branch”). Together, these categories provide a structured and fine-grained basis for evaluating compositional alignment.
@@ 3.1 Experimental Setting
Our analysis is based on T2I-CompBench++ [13], which provides curated prompts across attributes (color, shape, texture), spatial relations (2D and 3D), non-spatial relations, complex prompts, and numeracy. Each prompt is paired with images from multiple text-to-image models and annotated with human evaluation scores. All resources (prompts, images, and scores) come from the benchmark; our contribution is to analyze how evaluation metrics align with these annotations using outputs from SD v1.4, SD v2, Structured Diffusion [9], Composable Diffusion [21], Attend-and-Excite [2], and GORS [12].
We evaluate five embedding-based metrics (PickScore [16], CLIPScore [10], HPS [31], ImageReward [32], BLIP-2 [18]), two image-only metrics (CLIP-IQA [30], Aesthetic Score [28]), and five VQA-based metrics (B-VQA [12], DA Score [29], TIFA [11], DSG [5], VQA Score [20]), covering embedding similarity, perceptual quality, and VQA-style reasoning.
@@ 3.2 Correlation Analysis of Evaluation Metrics
CLIP [10] 0.282 0.291 0.535 0.369 0.439 0.276 0.315 0.223
PickScore [16] 0.263 0.270 0.516 0.299 0.432 0.167 0.139 0.337
HPS [31] 0.219 0.440 0.601 0.410 0.535 0.270 0.416 0.471
ImageReward [32] 0.580 0.520 0.734 0.394 0.512 0.424 0.401 0.484
BLIP2 [18] 0.250 0.287 0.546 0.369 0.353 0.235 0.416 0.366
Aesthetic [28] 0.056 0.195 0.078 0.136 0.061 0.051 0.123 0.036
CLIP-IQA [30] 0.092 0.078 -0.001 0.088 0.082 0.027 0.098 0.068
B-VQA [12] 0.610 0.388 0.690 0.255 0.371 0.372 0.330 0.444
DA Score [29] 0.772 0.463 0.711 0.318 0.453 0.488 0.297 0.462
TIFA [11] 0.684 0.336 0.423 0.311 0.351 0.519 0.195 0.526
DSG [5] 0.599 0.388 0.628 0.328 0.470 0.411 0.427 0.469
VQA Score [20] 0.678 0.405 0.701 0.533 0.495 0.638 0.339 0.473
@@
Several broader insights emerge from these results. First, no single metric achieves strong and consistent correlation across all compositional categories, indicating that reliance on a single signal is insufficient. Second, despite its widespread use [26, 23, 27, 1, 17, 14, 3, 24, 4, 19, 22], CLIP never ranks among the top metrics, underscoring its limitations as a standalone measure. Third, embedding-based metrics, particularly ImageReward and HPS, frequently appear among the strongest. Fourth, while VQA-based metrics are competitive, they are not uniformly superior and are occasionally outperformed by embedding-based approaches. Finally, image-only metrics such as CLIP-IQA and Aesthetic remain consistently weak, as expected since they do not assess text–image alignment.
@@ 3.3 Beyond Correlation: Regression Analysis of Metrics
CLIP [10] -0.290 0.000 -0.148 0.000 0.649 0.000 -0.444 0.000
PickScore [16] 2.467 0.000 1.564 0.000 0.000 0.000 -2.801 0.000
HPS [31] -0.097 0.761 0.475 1.143 0.629 0.000 2.048 1.277
ImageReward [32] 0.193 0.334 0.300 0.197 -0.095 0.223 0.008 0.053
BLIP2 [18] -0.420 0.000 0.264 0.000 0.594 0.000 0.510 0.000
Aesthetic [28] 0.229 0.000 0.260 0.000 0.023 0.000 -0.194 0.000
CLIP-IQA [30] -0.017 0.000 0.014 0.016 0.000 0.000 0.100 0.000
B-VQA [12] -0.092 -0.063 0.059 -0.264 0.035 0.000 0.071 0.059
DA Score [29] 0.754 0.132 0.236 0.428 0.000 0.000 -0.223 0.051
TIFA [11] 0.211 0.028 -0.015 0.003 0.061 0.214 -0.002 0.204
DSG [5] -0.043 0.068 0.101 0.000 0.023 0.000 0.161 -0.045
VQA Score [20] 0.112 0.082 0.064 0.319 0.068 0.246 -0.003 0.092
@@ Appendix A Correlation Study
CLIP [10] 0.275 0.276 0.530 0.345 0.651 0.349 0.306 0.210
PickScore [16] 0.227 0.263 0.492 0.314 0.558 0.205 0.175 0.376
HPS [31] 0.240 0.457 0.555 0.431 0.623 0.355 0.460 0.495
ImageReward [32] 0.585 0.555 0.737 0.403 0.543 0.496 0.405 0.482
BLIP2 [18] 0.259 0.333 0.541 0.344 0.620 0.299 0.412 0.364
Aesthetic [28] 0.078 0.195 0.068 0.152 0.095 0.077 0.097 0.061
CLIP-IQA [30] 0.114 0.080 0.019 0.070 0.015 0.023 0.108 0.018
B-VQA [12] 0.621 0.348 0.644 0.237 0.462 0.426 0.335 0.460
DA Score [29] 0.793 0.464 0.674 0.324 0.519 0.513 0.298 0.485
TIFA [11] 0.685 0.364 0.415 0.298 0.451 0.552 0.212 0.530
DSG [5] 0.569 0.392 0.598 0.332 0.507 0.355 0.441 0.450
VQA Score [20] 0.682 0.409 0.617 0.452 0.542 0.613 0.348 0.487
@@
CLIP [10] 0.208 0.211 0.392 0.287 0.347 0.201 0.224 0.154
PickScore [16] 0.193 0.192 0.373 0.229 0.341 0.122 0.100 0.241
HPS [31] 0.157 0.326 0.441 0.315 0.428 0.201 0.305 0.346
ImageReward [32] 0.434 0.388 0.549 0.310 0.408 0.313 0.294 0.349
BLIP2 [18] 0.179 0.203 0.389 0.286 0.280 0.170 0.303 0.264
Aesthetic [28] 0.039 0.138 0.054 0.104 0.047 0.037 0.083 0.026
CLIP-IQA [30] 0.065 0.055 -0.002 0.068 0.063 0.018 0.068 0.045
B-VQA [12] 0.456 0.279 0.512 0.195 0.293 0.267 0.231 0.322
DA Score [29] 0.603 0.337 0.534 0.247 0.357 0.364 0.206 0.347
TIFA [11] 0.559 0.246 0.329 0.266 0.292 0.405 0.155 0.400
DSG [5] 0.499 0.303 0.503 0.292 0.408 0.325 0.355 0.363
VQA Score [20] 0.512 0.292 0.516 0.422 0.390 0.481 0.243 0.352
@@
CLIP [10] 0
PickScore [16] 0
HPS [31] 4
ImageReward [32] 6
BLIP2 [18] 1
Aesthetic [28] 0
CLIP-IQA [30] 0
B-VQA [12] 0
DA Score [29] 4
TIFA [11] 3
DSG [5] 1
VQA Score [20] 6

### v2

# The First Workshop on Generative and Protective AI for Content Creation

@@ 1 Introduction
Recent advances in multi-modal generative models, such as Stable Diffusion [rombach2022high, podell2023sdxl, esser2024scaling] and DALL-E [ramesh2022hierarchical], have made it possible to generate high-quality, natural, and diverse images from textual descriptions at low cost and on a large scale. Alongside this progress, a parallel line of research has emerged around a central question: how can we faithfully assess the alignment between text and image, particularly in terms of compositional alignment?
Text–image compositional alignment refers to how well the objects, attributes, and relations described in a prompt are reflected in the generated image. A variety of metrics have been introduced to evaluate this alignment. Content-based metrics, such as VQAScore [lin2024evaluating], TIFA [hu2023tifa], DA Score [singh2023divide], DSG [cho2023davidsonian], and B-VQA [huang2023t2i], assess compositional properties through structured queries to the image. Embedding-based metrics, including CLIPScore [hessel2021clipscore], PickScore [kirstain2023pickscore], HPS [wu2023hps], and ImageReward [xu2024imagereward], measure alignment via text–image similarity in a shared representation space or by using models trained on human preference data. Image-only metrics, such as CLIP-IQA [wang2023exploring] and Aesthetic Score [schuhmann2022laion], instead focus on perceptual quality and visual appeal, complementing alignment-based evaluation. Together, these metrics determine how models are evaluated and compared, and much of the field’s reported progress is defined through their outcomes. Selecting appropriate metrics is therefore essential.
In addition to their role as evaluation tools, text–image alignment metrics have increasingly served as reward signals to enhance compositional generation in diffusion models through both inference-time and reinforcement learning strategies. For instance, ReNO [eyring2024reno] leverages multiple reward models, including CLIPScore [hessel2021clipscore], ImageReward [xu2024imagereward], PickScore [kirstain2023pickscore], and HPSv2 [wu2023hps], to guide gradient-based optimization of initial noise during inference, improving both faithfulness and aesthetics. In contrast, ImageSelect [karthik2023if] adopts a best-of-$N$ sampling strategy, generating multiple candidate images from different initial noises and selecting the one with the highest ImageReward [xu2024imagereward] score. On the training side, reinforcement learning methods like DPOK [fan2023dpok] fine-tune the diffusion model by maximizing ImageReward [xu2024imagereward], using online policy gradient to strengthen compositional alignment.
Despite their widespread use in both evaluation and generation, there has been no thorough and comparative analysis of how well these metrics reflect human judgment. To address this gap, this work presents a comprehensive evaluation of $12$ text–image compositional alignment metrics on $2400$ generated text–image samples spanning $8$ compositional categories [t2iplus], examining how closely each metric corresponds to human assessments.
@@ 2.1 Evaluation Metrics
Embedding-based metrics evaluate alignment by comparing text–image representations in a shared multimodal space or by leveraging models trained on human preferences. A common baseline is CLIPScore [hessel2021clipscore], which measures cosine similarity between CLIP embeddings. Preference-supervised variants include HPS [wu2023hps], which fine-tunes CLIP on human comparisons, and PickScore [kirstain2023pickscore], which learns from pairwise preference judgments. BLIP [li2023blip2] follows the embedding-similarity approach, comparing captions generated from images with the input text. Extending this idea, ImageReward [xu2024imagereward] adds a reward head trained on ranked human preference data, capturing both textual relevance and perceptual quality.
@@
VQA-based metrics assess compositional alignment by casting text–image consistency as a question answering task. Questions derived from the prompt are posed to a pretrained VQA model, with scores based on the correctness of its responses. VQAScore [lin2024evaluating] generates yes/no questions from the text, while TIFA [hu2023tifa] uses structured templates to cover objects, attributes, and relations. Variants target specific aspects: DA Score [singh2023divide] asks entity–attribute questions to test binding, DSG [cho2023davidsonian] converts the text into a scene graph to verify entities and relations, and B-VQA [huang2023t2i] decomposes the text into object–attribute pairs, querying each with BLIP-VQA and combining the probabilities.
@@
Image-only metrics assess perceptual quality independently of the prompt, providing complementary signals of realism and aesthetics. CLIP-IQA [wang2023exploring] predicts image quality by regressing CLIP embeddings against human quality annotations, while the Aesthetic Score [schuhmann2022laion] estimates aesthetic value from large-scale human ratings.
@@ 2.2 Benchmarks
Text–image compositional alignment benchmarks, such as T2I-CompBench++ [t2iplus], group prompts into categories that reflect major alignment challenges. These include: entity existence (all entities appear without omission or hallucination), attribute binding (objects match specified properties like color or shape), spatial relations (2D and 3D object arrangements, e.g., “a cube on a sphere” or “a ball inside a box”), non-spatial relations (functional interactions, e.g., “a man holding a guitar”), and numeracy (accurate object counts, e.g., “five birds on a branch”). Together, these categories provide a structured and fine-grained basis for evaluating compositional alignment.
@@ 3.1 Experimental Setting
Our analysis is based on T2I-CompBench++ [t2iplus], which provides curated prompts across attributes (color, shape, texture), spatial relations (2D and 3D), non-spatial relations, complex prompts, and numeracy. Each prompt is paired with images from multiple text-to-image models and annotated with human evaluation scores. All resources (prompts, images, and scores) come from the benchmark; our contribution is to analyze how evaluation metrics align with these annotations using outputs from SD v1.4, SD v2, Structured Diffusion [Feng2022TrainingFreeSD], Composable Diffusion [liu2022compositional], Attend-and-Excite [Chefer2023AttendandExciteAS], and GORS [huang2023t2i].
We evaluate five embedding-based metrics (PickScore [kirstain2023pickscore], CLIPScore [hessel2021clipscore], HPS [wu2023hps], ImageReward [xu2024imagereward], BLIP-2 [li2023blip2]), two image-only metrics (CLIP-IQA [wang2023exploring], Aesthetic Score [schuhmann2022laion]), and five VQA-based metrics (B-VQA [huang2023t2i], DA Score [singh2023divide], TIFA [hu2023tifa], DSG [cho2023davidsonian], VQA Score [lin2024evaluating]), covering embedding similarity, perceptual quality, and VQA-style reasoning.
@@ 3.2 Correlation Analysis of Evaluation Metrics
CLIP [hessel2021clipscore] 0.282 0.291 0.535 0.369 0.439 0.276 0.315 0.223
PickScore [kirstain2023pickscore] 0.263 0.270 0.516 0.299 0.432 0.167 0.139 0.337
HPS [wu2023hps] 0.219 0.440 0.601 0.410 0.535 0.270 0.416 0.471
ImageReward [xu2024imagereward] 0.580 0.520 0.734 0.394 0.512 0.424 0.401 0.484
BLIP2 [li2023blip2] 0.250 0.287 0.546 0.369 0.353 0.235 0.416 0.366
Aesthetic [schuhmann2022laion] 0.056 0.195 0.078 0.136 0.061 0.051 0.123 0.036
CLIP-IQA [wang2023exploring] 0.092 0.078 -0.001 0.088 0.082 0.027 0.098 0.068
B-VQA [huang2023t2i] 0.610 0.388 0.690 0.255 0.371 0.372 0.330 0.444
DA Score [singh2023divide] 0.772 0.463 0.711 0.318 0.453 0.488 0.297 0.462
TIFA [hu2023tifa] 0.684 0.336 0.423 0.311 0.351 0.519 0.195 0.526
DSG [cho2023davidsonian] 0.599 0.388 0.628 0.328 0.470 0.411 0.427 0.469
VQA Score [lin2024evaluating] 0.678 0.405 0.701 0.533 0.495 0.638 0.339 0.473
@@
Several broader insights emerge from these results. First, no single metric achieves strong and consistent correlation across all compositional categories, indicating that reliance on a single signal is insufficient. Second, despite its widespread use [rombach2022high, nichol2021glide, ruiz2023dreambooth, brooks2023instructpix2pix, kumari2023multi, kang2023scaling, chefer2023attend, podell2023sdxl, chen2023pixartalpha, li2023gligen, nguyen2024swiftbrush], CLIP never ranks among the top metrics, underscoring its limitations as a standalone measure. Third, embedding-based metrics, particularly ImageReward and HPS, frequently appear among the strongest. Fourth, while VQA-based metrics are competitive, they are not uniformly superior and are occasionally outperformed by embedding-based approaches. Finally, image-only metrics such as CLIP-IQA and Aesthetic remain consistently weak, as expected since they do not assess text–image alignment.
@@ 3.3 Beyond Correlation: Regression Analysis of Metrics
CLIP [hessel2021clipscore] -0.290 0.000 -0.148 0.000 0.649 0.000 -0.444 0.000
PickScore [kirstain2023pickscore] 2.467 0.000 1.564 0.000 0.000 0.000 -2.801 0.000
HPS [wu2023hps] -0.097 0.761 0.475 1.143 0.629 0.000 2.048 1.277
ImageReward [xu2024imagereward] 0.193 0.334 0.300 0.197 -0.095 0.223 0.008 0.053
BLIP2 [li2023blip2] -0.420 0.000 0.264 0.000 0.594 0.000 0.510 0.000
Aesthetic [schuhmann2022laion] 0.229 0.000 0.260 0.000 0.023 0.000 -0.194 0.000
CLIP-IQA [wang2023exploring] -0.017 0.000 0.014 0.016 0.000 0.000 0.100 0.000
B-VQA [huang2023t2i] -0.092 -0.063 0.059 -0.264 0.035 0.000 0.071 0.059
DA Score [singh2023divide] 0.754 0.132 0.236 0.428 0.000 0.000 -0.223 0.051
TIFA [hu2023tifa] 0.211 0.028 -0.015 0.003 0.061 0.214 -0.002 0.204
DSG [cho2023davidsonian] -0.043 0.068 0.101 0.000 0.023 0.000 0.161 -0.045
VQA Score [lin2024evaluating] 0.112 0.082 0.064 0.319 0.068 0.246 -0.003 0.092
@@ Appendix A Correlation Study
CLIP [hessel2021clipscore] 0.275 0.276 0.530 0.345 0.651 0.349 0.306 0.210
PickScore [kirstain2023pickscore] 0.227 0.263 0.492 0.314 0.558 0.205 0.175 0.376
HPS [wu2023hps] 0.240 0.457 0.555 0.431 0.623 0.355 0.460 0.495
ImageReward [xu2024imagereward] 0.585 0.555 0.737 0.403 0.543 0.496 0.405 0.482
BLIP2 [li2023blip2] 0.259 0.333 0.541 0.344 0.620 0.299 0.412 0.364
Aesthetic [schuhmann2022laion] 0.078 0.195 0.068 0.152 0.095 0.077 0.097 0.061
CLIP-IQA [wang2023exploring] 0.114 0.080 0.019 0.070 0.015 0.023 0.108 0.018
B-VQA [huang2023t2i] 0.621 0.348 0.644 0.237 0.462 0.426 0.335 0.460
DA Score [singh2023divide] 0.793 0.464 0.674 0.324 0.519 0.513 0.298 0.485
TIFA [hu2023tifa] 0.685 0.364 0.415 0.298 0.451 0.552 0.212 0.530
DSG [cho2023davidsonian] 0.569 0.392 0.598 0.332 0.507 0.355 0.441 0.450
VQA Score [lin2024evaluating] 0.682 0.409 0.617 0.452 0.542 0.613 0.348 0.487
@@
CLIP [hessel2021clipscore] 0.208 0.211 0.392 0.287 0.347 0.201 0.224 0.154
PickScore [kirstain2023pickscore] 0.193 0.192 0.373 0.229 0.341 0.122 0.100 0.241
HPS [wu2023hps] 0.157 0.326 0.441 0.315 0.428 0.201 0.305 0.346
ImageReward [xu2024imagereward] 0.434 0.388 0.549 0.310 0.408 0.313 0.294 0.349
BLIP2 [li2023blip2] 0.179 0.203 0.389 0.286 0.280 0.170 0.303 0.264
Aesthetic [schuhmann2022laion] 0.039 0.138 0.054 0.104 0.047 0.037 0.083 0.026
CLIP-IQA [wang2023exploring] 0.065 0.055 -0.002 0.068 0.063 0.018 0.068 0.045
B-VQA [huang2023t2i] 0.456 0.279 0.512 0.195 0.293 0.267 0.231 0.322
DA Score [singh2023divide] 0.603 0.337 0.534 0.247 0.357 0.364 0.206 0.347
TIFA [hu2023tifa] 0.559 0.246 0.329 0.266 0.292 0.405 0.155 0.400
DSG [cho2023davidsonian] 0.499 0.303 0.503 0.292 0.408 0.325 0.355 0.363
VQA Score [lin2024evaluating] 0.512 0.292 0.516 0.422 0.390 0.481 0.243 0.352
@@
CLIP [hessel2021clipscore] 0
PickScore [kirstain2023pickscore] 0
HPS [wu2023hps] 4
ImageReward [xu2024imagereward] 6
BLIP2 [li2023blip2] 1
Aesthetic [schuhmann2022laion] 0
CLIP-IQA [wang2023exploring] 0
B-VQA [huang2023t2i] 0
DA Score [singh2023divide] 4
TIFA [hu2023tifa] 3
DSG [cho2023davidsonian] 1
VQA Score [lin2024evaluating] 6